OpenStudy (anonymous):

Will give medal simplify the following expression 7x – 4y + 5z – 2x – 10y + z A. 9x – 14y + 6z B. 5x – 14y + 6z C. 5x – 6y + 6z D. 9x – 6y + 6z

OpenStudy (anonymous):

combine like terms sooo. 7x-2x= -4y-10y= 5z+z= and put them all together.

OpenStudy (anonymous):

sooo the answer is B.5x-14y+6z

OpenStudy (anonymous):

thank you so so much!!!!